# Empowering Autonomous Driving with Large Language Models: A Safety Perspective

### 18 Dec 2023 v3

@@ 4 LLM as Interactive Decision Maker: Interactive Planning by Behavior Prediction and State Machine
As with most existing works on LLM for AD, our previous case study focuses on such one-step planning or single-frame decision-making. We can further improve the performance and safety of LLM for driving tasks by explicitly considering the ego vehicle’s high-level behavior patterns (controlled by LLM) and the prediction and interaction with surrounding agents in multiple consecutive steps. Based on the first case study, the highlights of this second case study are as follows.
State Machine as Behavior Guidance for LLM Decision Maker: In practical driving situations, executing maneuvers like lane changes or unprotected left turns involves several steps. Vehicles must adhere to established rules or patterns to interact effectively with nearby traffic and accomplish these tasks. This necessitates that LLMs conform to specific behavioral constraints. We propose to integrate a state machine into the LLM through in-context learning or few-shot fine-tuning. This approach aims to ensure adherence to these essential regulations and established patterns while maintaining sufficient flexibility in decision-making.
@@
In Figure 5, we present our pipeline for interactive lane changing using LLM as the decision-maker. The framework is centered around the state machine which defines the basic behavior pattern of our LLM. The memory stores important past information about surrounding vehicles, helping the LLM make predictions of their intentions. The reflection module is to monitor the LLM and make sure the transition is valid from state to state and to give feedback to the LLM for in-context learning when the LLM violates transition constraints. The LLM determines transitions based on predefined rules and inferred information. The transition involves several checks:
@@
Safety Check: The LLM evaluates the possibility of collision if it takes certain actions transiting to the next state. In this particular example of interactive lane changing, the time-to-collision (TTC) is applied to ensure the proposed state won’t lead to a collision. The LLM will compare the TTC against a set threshold.
@@
Intention Prediction Module: Unlike the prediction for MPC, the intention prediction is to estimate the high-level behavior patterns of the surrounding vehicle, which doesn’t need to be very detailed but is important for interaction. We define the potential intention of surrounding agents as cooperative and aggressive. We use the time-to-collision (TTC) of surrounding vehicles to measure their aggressiveness levels. At every planning step, the LLM decision-maker will extract the surrounding vehicles’ TTCs with the past 3 steps and predict their corresponding intentions. We will give several human-labeled demonstrations when setting up the LLM.
Experimental Analysis: In the experiment, we demonstrate how the proposed framework enables the LLM to conduct interactive and safe motion planning in consecutive steps based on a predefined state machine. As shown in Figure 7 and Figure 6, we notice that the LLM can ensure safety in the behavior level by adhering to the three constraints. The LLM continually estimates the surrounding vehicle’s intention and the estimation is further confirmed during the proactive interaction state - "Attempt". After LLM makes its decision, the behavior safety checker and MPC will ensure the feasibility and safety in the behavior level and the control level, respectively.
@@ 5 Discussion: LLM as Other Roles for Safety
In previous case studies, we first naively leveraged LLM as a one-step behavior decision-maker and then further developed it into interactive planning with state machine guidance and behavior prediction for surrounding agents in multiple steps. Observing improved performance and safety in the case studies, we are going to discuss the possible usage of LLMs for other components in the AD software pipeline, as shown in Figure 1. We directly ask the question for ChatGPT-3.5 (e.g., prompt as “How can a large language model assist the perception module for safer autonomous driving?”) and summarize its responses below.
5.1 LLMs in Perception
@@
1) Multimodel Fusion: It is possible to consider multimode infusion with language input. By integrating information from both sensor data and language input, the perception module can create a more comprehensive understanding of the environment. This multimodal fusion enables the system to make more informed safer decisions by considering both visual information and contextual cues provided by natural language. 2) Semantic Object Recognition: LLMs can assist in recognizing and understanding objects in the environment based on their semantic context of safety. For instance, if a passenger says, “Watch out for the cyclist ahead”, LLMs can understand this information to prioritize and adapt the behavior accordingly, enhancing safety. 3) Adaptive Object Detection: LLMs can provide information that helps the perception module adapt its object detection algorithms based on specific scenarios. For example, if LLMs understand that the vehicle is in a construction zone, they can convey this information to the perception module, prompting the system to be more cautious and attentive to potential hazards. 4) Human-Interacted Perception: The language model can generate annotations or labels for objects detected by the perception module in the environment with natural language which could increase the user’s trust. Furthermore, LLMs can enhance the perception module’s ability to recognize and respond to human-related cues including the ones related to safety. 5) Reducing False Positives: The language model can help filter out false positives in object detection by considering contextual information. If a sensor generates a potential obstacle, the language model can verify its validity based on the overall context, reducing the likelihood of unnecessary interventions.
5.2 LLMs in Prediction
@@
1) Natural Language Inputs for Contextual Awareness: The language model in the prediction module can process natural language inputs (possibly from perception) to understand and infer the potential intentions of other drivers. For example, if the perception model or human user interprets "heavy traffic ahead," the prediction module with LLMs can understand it and adjust its expectations and predictions accordingly for safer operation. 2) Human-Centric Predictions: Language understanding can help the prediction module make more human-centric predictions by considering factors such as hand gestures, turn signals, or spoken commands from other drivers. This allows the autonomous vehicle to anticipate and respond to human behaviors more effectively, improving AV safety. 3) Behavioral Evaluation: The language model can assist in evaluating the driving behaviors and aggressiveness of surrounding cars by processing information such as traffic reports or user-generated input. This helps the prediction module adjust its predictions based on the perceived driving styles of other vehicles. 4) Learning from Historical Data: Due to the memorization ability, the language model can facilitate learning from historical data by interpreting user feedback or incorporating information from past interactions. This adaptive learning approach allows the prediction module to improve its predictions over time based on real-world experiences.
5.3 LLMs in Simulation
@@
1) User Specific Scenario Generation and Variation: The language model can generate natural language descriptions of diverse driving scenarios by user input for safety concerns, allowing the simulation module to create a wide range of realistic and challenging situations for testing and training in a safety perspective. This helps in ensuring that the autonomous system is well-prepared for various real-world conditions. 2) Human-Like Interaction: The language model can simulate human-like interactions by generating realistic communication between simulated drivers, pedestrians, and other entities. This enhances the realism of the simulation, allowing the autonomous system to practice responding to natural language cues and gestures for safety purposes. 3) Simulation Annotation and Analysis: The language model can assist in annotating simulation data by generating descriptions or labels for different events and entities, which further the AV development. 4) Fault and Unsafe Edge Case Analysis: The language model can aid in summarizing and generating natural language descriptions of simulated faults and unsafe edge cases in the simulation environment. This helps the developers to better understand and improve the AD systems.
@@ 6 Conclusion
In conclusion, our presented framework explores the integration of an LLM in an autonomous driving system, fortified by a safety verifier within the software pipeline to facilitate in-context safety learning. The versatility of LLMs is leveraged across perception, prediction, and planning, thereby augmenting overall system safety, as extensively elaborated in this paper. Through two compelling case studies, we demonstrate the efficacy of employing LLMs as behavior decision-makers, showcasing notable enhancements in both performance and safety. We intend to broaden the safety perspective within the research community concerning the utilization of LLMs in the domain of autonomous driving. We believe that this paper catalyzes further exploration and innovation in ensuring the safety, reliability, and interpretability of autonomous vehicles.

### 22 Mar 2024 v4

@@ 4 LLM as Interactive Decision Maker: Interactive Planning by Behavior Prediction and State Machine
As with most existing works on LLM for AD, our previous case study focuses on one-step planning or single-frame decision-making. We can further improve the performance and safety of LLM for driving tasks by explicitly considering the ego vehicle’s high-level behavior transitions and the interaction with surrounding agents in multiple consecutive steps. In Figure 5, besides the MPC verifier we proposed previously, we further design the state machine framework as behavior transition guidance, the memory module for intention prediction, and the reflection module for behavior-level safety checks and in-context learning. We will explain them in detail in the following.
State Machine as Behavior Guidance for LLM Decision Maker: In practical driving situations, executing maneuvers like lane changes or unprotected left turns involves several steps. Vehicles must adhere to established rules or patterns to interact effectively with nearby traffic and accomplish these tasks. This necessitates that LLMs conform to specific behavioral constraints. We propose to integrate a state machine into the LLM through in-context learning. This approach aims to ensure adherence to these essential regulations and established patterns while maintaining sufficient flexibility in decision-making.
@@
In Figure 5, we present our pipeline for interactive lane changing using LLM as the decision-maker. The framework is centered around the state machine which defines the basic behavior pattern of our LLM. The memory stores important past information about surrounding vehicles, helping the LLM make predictions of their intentions. The reflection module is to monitor the LLM and make sure the transition is valid from state to state and to give feedback to the LLM for in-context learning when the LLM violates hard transition constraints. The LLM determines transitions based on predefined rules and inferred information. The transition involves several checks:
@@
Safety Check: The LLM evaluates the possibility of collision if it takes certain actions transiting to the next state. In this particular lane-changing example, the time-to-collision (TTC) is applied to ensure the proposed state won’t lead to a collision. The LLM will compare the TTC against a set threshold.
@@
Intention Prediction Module: Unlike the prediction for MPC, the intention prediction is to estimate the high-level behavior patterns of the surrounding vehicle, which doesn’t need to be very detailed but is important for interaction. We define the potential intention of surrounding agents as cooperative and aggressive. We use the time-to-collision (TTC) of surrounding vehicles as input to the LLM for prediction. At every planning step, the LLM decision-maker will extract the surrounding vehicles’ TTCs with the past 3 steps and predict their corresponding intentions. We give several human-labeled demonstrations when setting up the LLM.
Collision Rate | Success Rate | |
DriveLikeAHuman | 47.1% | 41.2% |
Ours | 0 | 100% |
Ours w/o failsafe | 23.1% | 76.9% |
Ours w/o reflection | 0 | 92.3% |
Experimental Analysis: In this study, we evaluate our proposed framework using the HighwayEnv simulation platform. As depicted in Figure 6, our framework successfully guides the LLM to perform safe motion planning in sequential steps, relying on a state machine, along with prediction and reflection modules. Figure 7 visualizes the lane-changing scenario, showing the LLM’s continuous reasoning and interaction with nearby vehicles under complex conditions. This figure also details the state transitions within the decision-making process. we compared our approach with the open-source DriveLikeAHuman Fu et al. (2024) framework in terms of safety (collision rate) and the success rate of lane changes. The findings, presented in Table 2, indicate a significantly higher rate of collisions and aborts with the DriveLikeAHuman’s naive chain-of-thoughts strategy. In contrast, our method not only ensures safety but also exhibits a remarkable success rate in a variety of generated scenarios, highlighting the efficacy and generalizability of our bi-level interactive planning framework. The final two columns of Table 2 showcase the significance of our framework’s components through an ablation study.
@@ 5 Discussion: LLM as Other Roles for Safety
We discuss the possible usage of LLMs for other components in the AD software pipeline, as shown in Figure 1. We directly ask ChatGPT-3.5 (e.g., prompt as “How can a large language model assist the perception module for safer autonomous driving?”) and summarize its responses below.
@@
LLMs for Perception. 1) Multimodal Fusion: It is possible to consider multimodal infusion with language input. By integrating information from both sensor data and language input, the perception module can create a more comprehensive understanding of the environment. This multimodal fusion enables the system to make more informed safer decisions by considering both visual information and contextual cues provided by natural language. 2) Semantic Object Recognition: LLMs can assist in recognizing and understanding objects in the environment based on their semantic context of safety. For instance, if a passenger says, “Watch out for the cyclist ahead”, LLMs can understand this information to prioritize and adapt the behavior accordingly, enhancing safety. 3) Adaptive Object Detection: LLMs can provide information that helps the perception module adapt its object detection algorithms based on specific scenarios. For example, if LLMs understand that the vehicle is in a construction zone, they can convey this information to the perception module, prompting the system to be more cautious and attentive to potential hazards.
@@
LLMs for Prediction. 1) Natural Language Inputs for Contextual Awareness: The language model in the prediction module can process natural language inputs (possibly from perception) to understand and infer the potential intentions of other drivers. For example, if the perception model or human user interprets "heavy traffic ahead," the prediction module with LLMs can understand it and adjust its expectations and predictions accordingly for safer operation. 2) Human-Centric Predictions: Language understanding can help the prediction module make more human-centric predictions by considering factors such as hand gestures, turn signals, or spoken commands from other drivers. This allows the autonomous vehicle to anticipate and respond to human behaviors more effectively, improving AV safety. 3) Behavioral Evaluation: The language model can assist in evaluating the driving behaviors and aggressiveness of surrounding cars. This helps the prediction module adjust its predictions based on the perceived driving styles of other vehicles.
@@
LLMs for Simulation. 1) User Specific Scenario Generation and Variation: The language model can generate natural language descriptions of diverse driving scenarios by user input for safety concerns, allowing the simulation module to create a wide range of realistic and challenging situations for testing and training in a safety perspective. This helps in ensuring that the autonomous system is well-prepared for various real-world conditions. 2) Human-Like Interaction: The language model can simulate human-like interactions by generating realistic communication between simulated drivers, pedestrians, and other entities. This enhances the realism of the simulation, allowing the autonomous system to practice responding to natural language cues and gestures for safety purposes. 3) Simulation Annotation and Analysis: The language model can assist in annotating simulation data by generating descriptions or labels for different events and entities, which further the AV development.
@@ 6 Conclusion
In conclusion, our presented framework explores the integration of an LLM as an intelligent decision-maker for autonomous driving, fortified by a safety verifier feedback for in-context safety learning. Through two case studies, we demonstrate the efficacy of our approach, showcasing notable enhancements in both performance and safety. We further discuss the potential usage of the LLM for other components. This paper intends to broaden the safety perspective within the autonomous driving community concerning the utilization of LLMs. The future directions and remaining challenges include testing this framework in real-world environment and handling ambiguity, biases, and inconsistencies in LLM outputs.